OpenStudy (anonymous):

if the centre of a circle is (-6,8) and it passes through the origin , then equation of it's tangent at the origin is a)2y=x b)4y=3x c)3y=4x d)3x+4y=0

OpenStudy (anonymous):

@mathstudent55 @mathmath333

Parth (parthkohli):

Hint: the tangent is perpendicular to the radius. Find the equation of the radius and then find the equation of the line perpendicular to it.

OpenStudy (anonymous):

equation of tangent is 4x+3y=0. How to find the equation of line perpendicular to it ?

OpenStudy (anonymous):

is it 4y+3x=0 ?

Parth (parthkohli):

Somewhat. What would the slope of the perpendicular line be? Also, you know that the perpendicular line/tangent passes through (0,0).

OpenStudy (anonymous):

the slope of the tangent is 3/4 so tht of it's perpendicular should be -4/3 and equation of line passing through (0,0) with slope -4/3 is 4y+3x=0